OpenStudy (anonymous):

what is the slope of 2x=5y+30?

OpenStudy (nincompoop):

arrange the equation into slope-intercept y = mx + b

OpenStudy (anonymous):

it is a line equation. Isolate y and your a value (y=ax+b) will be your slope

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

But when I do, I get 5y=2x-30? Now what... lol

OpenStudy (nincompoop):

are you sure it is 2x?

OpenStudy (anonymous):

slope is -2/5 and y int = (0,-6)

OpenStudy (anonymous):

isolate y. So:\[y = \frac{ 2 }{ 5 }x + \frac{ 30 }{ 5 }\] You can make 30/5 equal to 6 and your slope will be the a value of (y=ax+b), so a is obviously equal to 2/5 :)

OpenStudy (nincompoop):

we are not suppose to give out answers

OpenStudy (anonymous):

sorry typo: -6

OpenStudy (anonymous):

okay, well sorry then

OpenStudy (anonymous):

well what do u want me to say then if you're all mad .kk wtvr good luck on your math hailee